# Lessons from pendulums: A design comparison of three lab activities

### v1

@@ IV.3 Tufts
At Tufts, the design of the pendulum lab aims to create an environment in which students recognize the need for experimental practices. An overarching goal for the course, which this first lab activity seeks to make clear to students, is for students to direct their own activity, make decisions, and take actions that are reasonable to them. This perspective views student agency as opposite to the “doing school” framing, where students’ activity is prescribed by instructions and is about arriving at the predetermined correct answer. In “doing school” students distance themselves intellectually from what is taking place: what they do matters to the extent that the instructor determines it to be correct.
To promote this particular vision of student agency (or, in a negative formulation, to avoid students’ taking up “doing school”), the overall amount and detail of the instructional guidance is significantly scaled back. This for the first two lab sessions, the instruction sheets are half a page and that is the extent of the formal instructions for the activity. In other words, the design shifts onto students responsibility for figuring out how to approach, organize, and carry out their experimentation and for generating knowledge claims from that experimentation.
In addition to shifting responsibility onto students, the minimal direction from the instructions can promote in students a genuine feeling of being in control. One hope is that students feel empowered by the agency and responsibility afforded to them. At the same time, students may feel discomfort with this responsibility and not knowing what to do: learning to manage these feelings and to develop positive affective responses to not-knowing is a desired outcome of the lab. In other work, we have referred to this type of affective development as meta-affective learning [jeon_2023, radoff_its_2019]).
In the first lab activity, the written instructions provide a challenge for students: “How precisely can you measure the period of a pendulum?” The instructions further clarify that students are responsible for quantifying the precision of their measurements and for explaining their experimental technique. Students have access to a large assortment of materials they can use to construct pendulums. Although these materials are pre-selected by the lab coordinator, the variety of materials affords countless different ways to build a pendulum. The session includes time for students to share their techniques and results.
Throughout the building of the pendulum and data collection, the teaching assistants ideally focus on listening and supporting the substance of student ideas. For example, many students turn to something like standard deviation to quantify uncertainty, and, in these cases, TAs will typically respond along the lines of “What does standard deviation mean to you?” or “What does it tell us about data?” If (and when) students are unclear about the meaning of standard deviation – if it seems like standard deviation is a black box algorithm to produce correct answers – the TA should tell them to focus on approaches that make sense to them. There is no instruction or guidance toward established statistical measures of variation. In concert with the minimal formal structure, the nature and tone of TA interactions emphasize the substance of student ideas. In this, and throughout the course, Hammer and the TAs emphasize that it is respectable to be unsure, and that that in itself is worth explaining.
In the second week, the instruction sheet prompts students to use their technique to test the claims of Galileo. Specifically, the instruction sheet says:
@@
This is the spring 2024 version of the instruction sheet. Every year Hammer solicits input from the TAs to edit the wording, which allows the instructions to be responsive to student ideas and builds up the agency of the TAs themselves. Indeed, the first sentence above, about student conjectures, was written in response to actual discussions in the class. A year-over-year consistent feature of this experimental prompt is the mention of Galileo’s claims.
Semantically, the question is not confirmatoryand the following paragraph begins “Please take care to be honest about your data and your findings!” Still, the inclusion of Galileo’s name in the instructional prompt often leads students to expect to confirm his claims (the famous scientist must be correct!). Although it cues expectations to confirm some external authority, potentially tipping students into a “doing school” mode, it also sets up an empirical discrepancy. In most sections, some lab groups will produce data that indicates (to them and/or to a knowing instructor) a potential relationship between amplitude and period. This potential discrepancy generates a need for experimental practices; data visualization, uncertainty estimation, and an understanding of the production of data have functional relevance to the genuine need to make sense of this discrepancy.
In the third session, lab groups present and discuss their conclusions. The whole class discussion is termed a “meaning-making session,” where the whole class attempts to come to a consensus regarding the experimental question. The instruction sheet for this session includes a translated quote from the 11th century Arabic/Persian scientist al-Hasan ibn al-Haytham, describing how seekers of truth should only submit to argument and demonstration and “make himself an enemy of all that he reads.” The instruction sheet explicitly says that the values of reasoning and empirical evidence, as opposed to deference to authority, are the core values of this course. In addition to presenting their work to their peers, students discuss and reflect on this quote in light of their work in the first two weeks.
Throughout the three-week activity, the instructional team emphasizes that they do not grade for the correctness of students’ outcome but for the quality of their engagement in conducting their investigation (including their professionalism) When the labs are complete, with comments and scores ready for students, Hammer brings up the topic in lecture, first to affirm the independence with respect to mass (and to discuss how that coheres with Newtonian theory), and second to consider the variation with amplitude. Almost always, students’ data provides the basis for claiming there is a dependence, and that sparks a discussion about the role of authority in science.
@@ V Comparisons
We all begin the semester with a pendulum lab we design (and redesign) based on a resources-based view of student knowledge, reasoning, and framing, with shared objectives regarding epistemological framing: Labs are where students can and should engage in constructing knowledge through empirical study of physical phenomena. Put another way, we emphasize to students that the goal is not “getting the right answer,” as they might expect, but rather learning to engage in high-quality, intellectually rigorous experimentation. But we design these labs in different ways. Table 1 lists the differences across our curricula that we identified in this work.
@@
In this section, we compare our approaches with respect to presentation, written guidance, and responsive teaching. We selected these particular aspects of the curricula to compare because they are all tightly connected to our main theoretical commitment to student agency. For each of us and for each of these aspects, these design decisions were informed by framing and agency. Thus, to make sense of the differences in our curricula, we must focus on the design reasoning involved in these decisions: the ideas, intuitions, and conjectures we considered in thinking through questions of how, why, and to what end and in responding to context-dependent constraints.
@@ V.1 Presentation of the task
At Cornell and Tufts, the labs ask students to evaluate a given model that has the period of the pendulum independent of amplitude and mass; at UWB, the prompt is neutral, asking that students investigate the question: Does the period of a pendulum depend on the amplitude of the swing? At Cornell, the instructions present an equation for the period of a pendulum, described as part of the simple harmonic model along with some explanation of what a model is. At Tufts, the instructions present Galileo’s claims that the period does not depend on amplitude or mass. Additionally, at Cornell, the lab unit as a whole is introduced as a model-testing one from the outset while at Tufts it is only in the second (of three) sessions that Galileo’s claims are introduced, generally with reference to student conjectures or findings from the first session.
Part of the reasoning for providing a model that students are meant to test, as we do at Cornell and Tufts, is to cue up expectations to confirm the given model. For Hammer, the mention of Galileo sets up an eventual reflective discussion about the role of authority in science and the importance of evidence and argumentation. For Holmes, the similar approach with a mathematical model sets up an epistemological version of “productive failure” [kapur_designing_2012]. The expected discrepancy between the provided model and students’ data generates an epistemological rich context for underscoring the importance of empirical model-testing. One reason for the more granular in-lab instructions at Cornell is to guide students to find the discrepancy.
@@
These different endpoints are variations on a theme. That is, it is appropriate to say that we all agree on the important epistemological expectations we aim to communicate to students in the first lab. At Cornell and Tufts, the elicit-confront-resolve approach serves to underscore the emphasis on the substance and quality of students’ work; at UWB, this emphasis is achieved by keeping the phrasing neutral: from the outset the focus is always on students’ data and students’ work. These variations in how we present the task mainly reflect our different perceptions of the student populations in our courses. They may also reflect differences in personal preferences: model-testing is a central and specific idea for the designers at Cornell while at Tufts and UWB the focus is on constructing scientific knowledge and “science as a refinement of everyday thinking.”
@@ V.2 Written guidance
One of the main differences among our curricula – indeed the initial inspiration for this paper – is in the written guidance we provide (or do not provide) for students to accomplish the task. The lab instructions at Cornell and UWB specify particular tasks and activities students should do. The lab instructions at Tufts specify only features of the goal for the day.
@@
For Holmes and Scherr, their relatively more detailed instructions address a concern that students may feel lost or overwhelmed to a degree that would inhibit their learning. The concern for Hammer, on the other hand, is that the length and specifications of instructions can cue students’ “doing school” framing. They may use t’, for example, without understanding, following instructions provided by authority. If there are no instructions to parse, the possibility of the instructions cuing or stabilizing that framing is not possible.
@@
Holmes and Scherr take the perspective that this common framing can be a potentially productive resource: Use student expectations that doing lab means following instructions, and design those instructions carefully to direct them towards productive activity. For example, they direct students to make a histogram, but leave the details to them. In subsequent sessions, instructions become sparser, giving students more responsibility for their decisions.
@@
Having the students use t’ to quantify distinguishability plays out differently at UWB and Cornell. At UWB, it guides students to consider the possibility of a subtle dependence. At Cornell, finding that dependence conflicts with the model they likely expected to confirm. At both institutions, it is up to the students to decide what to conclude. In this way, both curricula uses students’ willingness to “do school” as a way to build towards and cultivate agency. By guiding students towards productive activities, such as iterating and refining their experimental apparatus, the instructions ensure that all students engage in these tasks.
@@
Here, we see a difference in our theoretical stances towards student behavior and learning, albeit a fine-grained and nuanced one. An overarching goal of our labs is to shift how students frame their work, to be less about “doing school” and more about doing (experimental) physics. Yet, we have differing beliefs about the degree to which “doing school” inhibits productive scientific work and learning, and about the consequences of disrupting this framing. We also have different perspectives on the risks of students’ feeling lost or frustrated: Scherr and Holmes view these emotions as a barrier to engagement and participation, whereas Hammer hopes that by wrestling with such feelings students can grow from them, that the lab is a first opportunity for meta-affective learning[radoff_its_2019]. Important across all labs is that the instructors help students feel safe to try things and explore, including to trust that their grades are based on arrive at prescribed outcomes.
@@
Recent work in these labs has highlighted the complexity of the epistemic and affective dynamics of students’ engagement. In the Cornell labs, Sundstrom et al. [sundstrom_instructing_2023] observed that successful instructor interventions to shift students away from confirmation framing were not always associated with an increase in measures of student agency. On the other hand, in the Tufts labs, Descamps et al. [descamps_dynamics_2024], observed students clearly in a confirmation framing (assuming that a possible trend in their data was caused by experimental error) and yet, surprisingly, also taking up epistemic agency. Additionally, Jeon et al. [jeon_2024] highlight a students’ complex affective experiences of confusion in the Tufts labs: in conversation with her lab mates she insisted that she “hates physics” because their confusing experimental results defied her expectations that physics should align with common sense. At the same time, this students’ feelings of confusion motivated her to take up more intellectual agency during the lab. While thinking in terms of agentive and affective tradeoffs might be a useful instructional intuition, at least within non-traditional labs, it is not so clear-cut as “tradeoff” might imply.
@@
As theoretical constructs, framing and agency provide insight into various kinds of design decisions, but they can be taken up in ways that lead to different outcomes. One might focus on decision-making agency or on framing as a phenomenon of an individual’s cognition, but these are far from the only way to understand and apply framing and agency to pedagogy. Additionally, we cannot ignore the influence of secondary and contextual factors in the divergence of our in-lab guidance structures. At Cornell and UWB, that the instructions go beyond overviewing the main task of the day allows them to work in “mini-activities” to build up collaboration skills. The minimal lab instructions at Tufts have the effect of shifting various aspects of classroom management onto the TAs, which is aligned with Hammer’s active interest in TA education, which we take up in the next section.
@@ V.3 Responsive Teaching
We all encourage responsiveness, but, with minimal written guidance, the Tufts labs depend more critically on the instructors’ practices of attending, interpreting, and responding to their students’ thinking. In part, the differences in our expectations of instructors reflect differences in expectations for the materials. Thinking Critically in Physics Labs, the core of the materials used at Cornell and UWB, is designed for wide dissemination and use. It is difficult in general to rely on instructor responsiveness, and so these materials provide more detailed in-lab instructions to guide students toward the intended, rich arenas of experimental inquiry: making useful plots, quantifying uncertainty and refining one’s experimental method, generating conclusions and sifting through empirical conflicts.
The Tufts labs, in contrast, are only for use at Tufts. That has the advantages and disadvantages of the instructors’ freedom and responsibility to use their judgment as they work with students. We see this as a second “agentive tradeoff” that comes with the more detailed in-lab instructions, analogous to the tradeoff for students: The instructors, too, are guided and constrained with respect to what they may notice and how they may respond.
At UWB, Scherr expects the instructors, who are all faculty, to act with autonomy and agency, including to modify the materials and to notice, assess, and respond to student work during implementation. Scherr’s own modifications to the Thinking Critically in Physics Labs curriculum also reflects this aspect of her context: some of the material addressed in the homework tutorials at Cornell are primarily covered in the instructor mini-lectures at the beginning of each class. When Scherr introduces the activities around quantitatively comparing datasets, she uses examples from her personal and research interests; other instructors use different examples. Still, the detailed in-lab instructions function to provide clear representation of pedagogical intentions, while at the same time taking a familiar form, and in these ways may facilitate instructors’ adoption.
@@
At Tufts in contrast, part of the objective for Hammer is TA education and preparation. The TAs are a mix of graduate students and undergraduates, the latter often coming to the position after successful experiences as Learning Assistants [Otero_learningassistants] in the highly interactive lectures. The minimal in-lab instructions have the effect of shifting responsibility to the TAs for noticing, interpreting, and responding to students. Often this involves helping the students to shift in their framing of lab, such as when students look to the TA for prescriptive guidance over how to proceed.
Hammer’s practice has been to meet with the team of lab TAs for several hours each Friday. The TAs recount snippets of what they noticed in their students’ work over the week, and they choose examples of students’ written work. They compare interpretations and ideas for they might respond. They also discuss plans for coming week, including to suggest refinements to the coming assignment’s wording and design. Experience at Tufts suggests these meetings play an important role in ensuring TAs implement the curriculum as intended and develop skills in Responsive Teaching.

### v2

@@ IV.3 Tufts
At Tufts, the written instructions change in small ways from semester to semester; examples are attached as supplemental materials. They are designed to create an environment in which students recognize the need for experimental practices, in particular of attention to precision and of questioning authority. This fits an overarching goal for the course, including the "content" focused lecture portion, that students learn how to learn, including to assess the quality of their knowledge, to make decisions, and to take actions they see as reasonable. This perspective views students, especially at the start of the semester, as poised to frame lab as following prescribed steps toward predetermined correct results, "doing school" [46] in ways most have experienced. The labs try to promote "doing science," to shift responsibility onto students for figuring out how to approach, organize, and carry out their experimentation and for generating knowledge claims from that experimentation. To that end, the written materials are minimal, just half a page long for each of the first two sessions. They present the empirical challenge but provide no instructions on how to accomplish it.
The hope is that students feel empowered by the agency and responsibility afforded to them. Part of that, instructors anticipate, is students’ learning to manage feelings of uncertainty over what to do, without the clarity of steps to follow. Ideally, they come to enjoy the autonomy, including the challenges of not-knowing. That is, the labs are designed and taught in ways that may promote "meta-affective learning" [44, 62]). In labs as in the courses as a whole, instructors and teaching assistants try to recognize confusion as productive, responding to students favorably (e.g. "great!") when they express feeling unsure and prompting them to elaborate (e.g., "talk about what has you confused").
The first activity asks students: “How precisely can you measure the period of a pendulum?” The instructions clarify the answer should be quantitative, "e.g. to within 0.2 seconds, or 0.1, 0.0000001, whatever you can manage." The TA shows students materials available; there is a rich assortment. From there, they are responsible for building their own pendulums and deciding both how to be precise and how to quantify that precision. The session includes time for students to share their techniques and results.
Throughout the building of the pendulum and data collection, the TAs ideally focus on listening and supporting the substance of student ideas. Often students turn to standard deviation to quantify uncertainty, based on their experience in labs, and, in these cases, TAs respond along the lines of “What does standard deviation mean to you?” or “What does it tell us about data?” Most students are unclear about what it means; they see it as a black box algorithm they are "supposed to" use, a part of framing lab as "doing school." The TA should discourage using algorithms in this way; students should approach labs in ways that make sense to them.
With limited time, and to avoid prompting students’ expectations of following prescribed steps, there is no instruction or guidance toward established statistical measures of variation or representational forms. A tacit element of the reasoning for this choice, which this comparative analysis brought to light, is a belief that students have sufficient capabilities with statistics and data representation to generate sensible approaches. If there were more time available, it would be possible to follow students’ inventions with explicit instruction, as in Thinking Critically in Physics.
The instructions for the second week present the task as follows:
@@
This is the spring 2024 version of the instruction sheet. Every year, the lab teaching team reads and considers possible revisions. Indeed, the first sentence above, about student conjectures, was written in response to discussions TAs reported having in their sections. This allows the instructions to be responsive to student ideas and builds up the agency of the TAs themselves [29]. A year-over-year consistent feature of this experimental prompt is the mention of Galileo’s claims.
Semantically, the question is not confirmatory and the following paragraph begins “Please take care to be honest about your data and your findings!” (That sentence was added based on observations in previous semesters, of students setting their own evidence aside.) Still, the inclusion of Galileo’s name in the instructional prompt typically leads students to expect to confirm his claims (the famous scientist must be correct!). That expectation sets up an empirical discrepancy. In most sections, some lab groups will produce data that indicates (to them and/or to a knowing instructor) that period does change with period. This potential discrepancy generates a need for experimental practices; data visualization, uncertainty estimation, and an understanding of the production of data have functional relevance to the genuine need to make sense of this discrepancy. [^5] The expectation to confirm Galileo may also reflect or cue students’ framing lab as doing school, but see [14] for evidence it can support productive experimental practices.
In the third session, lab groups present and discuss their conclusions. [^6] In 2026, as we prepare this manuscript, the lab team decided to keep the pendulum lab to two sessions. This discussion is termed a “meaning-making session,” where the whole class attempts to come to a consensus regarding the experimental question. The instruction sheet for this session includes a translated quote from the 11th century Arab scientist and mathematician al-Hasan ibn al-Haytham, describing how seekers of truth should only submit to argument and demonstration and “make himself an enemy of all that he reads” [68]. The instruction sheet explicitly says that the values of reasoning and empirical evidence, as opposed to deference to authority, are core to the course. In addition to presenting their work to their peers, students discuss and reflect on this quote in light of their work in the first two weeks.
Throughout the three-week activity, the instructional team emphasizes that they do not grade for the correctness of students’ results but for the quality of their engagement in conducting their investigation. When the labs are complete, with comments and scores ready for students, the instructor brings up the topic in the main course lecture, first to affirm the independence with respect to mass (and to discuss how that coheres with Newtonian theory), and second to consider the variation with amplitude. Almost always, students’ data provides the basis for claiming there is a dependence, and that sparks a discussion about the role of authority in science.
@@ V Comparisons
We all begin the semester with a pendulum lab we design (and redesign) based on a resources-based view of student knowledge, reasoning, and framing, with shared objectives regarding epistemological framing: labs are where students can and should engage in constructing knowledge through empirical study of physical phenomena. We all emphasize to students that the goal is not “getting the right answer,” as they might expect, but rather learning to engage in high-quality, intellectually rigorous experimentation. But we design these labs in different ways. Table 1 lists the differences across our curricula that we identified in this work.
@@
In this section, we compare our approaches with respect to presentation, written guidance, and responsive teaching. We selected these aspects of the curricula to compare because they connect tightly to our goals of student agency and shared theoretical commitments regarding framing and agency. To make sense of the differences in our curricula, we focus on the design reasoning involved in our decisions.
@@ V.1 Presentation of the task
At Cornell, the instructions present an equation for the period, described as part of the simple harmonic model, along with some explanation of what a model is. At Tufts, the instructions present Galileo’s claims that the period does not depend on amplitude or mass. Additionally, at Cornell, the lab unit as a whole is introduced as a model-testing one from the outset while at Tufts it is only in the second (of three) sessions that Galileo’s claims are introduced, with reference to student ideas or findings from the first session. At UWB, the prompt is neutral, simply asking students to investigate whether the period of a pendulum depends on the amplitude.
Part of the reasoning for providing a model that students are meant to test, as at Cornell and Tufts, is to cue up expectations of confirmation. For Hammer, the mention of Galileo sets up an eventual reflective discussion about the role of authority in science and the importance of evidence and argumentation. For Holmes, the similar approach with a mathematical model sets up an epistemological version of “productive failure” [48]. The expected discrepancy between the provided model and students’ data generates an epistemological rich context for underscoring the importance of empirical model-testing. One reason for the more granular in-lab instructions at Cornell is to guide students to find the discrepancy.
@@
We all agree on the important epistemological expectations we aim to communicate to students in the first lab. We all present the task, with an elicit-confront-resolve approach or neutral phrasing, as a way to support students’ framing the lab as centered on their reasoning. The differences reflect our perceptions of the student populations in our courses. They also reflect differences in personal preferences: model-testing is a central and specific idea for the designers at Cornell while at Tufts and UWB the focus is on constructing scientific knowledge and “science as a refinement of everyday thinking.”
@@ V.2 Written guidance
One of the main differences among our curricula – indeed the initial inspiration for this paper – is in the written guidance we provide (or do not provide) for students. The lab instructions at Cornell and UWB specify particular tasks and activities; the instructions at Tufts do not.
@@
For Holmes and Scherr, their more detailed instructions address a concern that students may feel lost or overwhelmed to a degree that would inhibit their learning. The concern for Hammer, on the other hand, is that the length and specifications of instructions can cue students’ “doing school” framing. If there are no instructions to parse, the possibility of the instructions cuing or stabilizing that framing is not possible.
@@
Holmes and Scherr take the perspective that this common framing can be a potentially productive resource: use student expectations that doing lab means following instructions, and design those instructions carefully to direct them towards productive activity, such as to make a histogram. In subsequent sessions, instructions become sparser, giving students more responsibility for their decisions.
@@
Having the students use t’ to quantify distinguishability plays out differently at UWB and Cornell. At UWB, the instructions guide students to consider the possibility of a subtle dependence. At Cornell, finding that dependence conflicts with the model they likely expected to confirm. At both institutions, it is up to the students to decide what to conclude. In this way, both curricula use students’ willingness to “do school” as a way to scaffold productive activities, reducing the guidance over time (within the first lab activity and across the term as a whole) to cultivate agency.
@@
Here, we see a difference in our theoretical stances towards student behavior and learning, within the same general framework. We differ in our expectations about the dynamics of students’ framing, in particular over when and how they may come to frame lab as "doing science" for themselves. Can students’ willingness to adhere to instructions —as "jumping through hoops" [57] or "doing school" [46]—support their progress toward productive scientific work, or is it something to disrupt from the start? We also have different perspectives on the risks of students’ feeling lost or frustrated: Scherr and Holmes view these emotions as a barrier to engagement and participation, whereas Hammer hopes that students grow from wrestling with such feelings, that the lab is an opportunity for meta-affective learning [62]. (We all, meanwhile, agree it is important that students feel safe to try things and explore, including to trust that their grades are not based on arriving at prescribed outcomes.)
Recent work in these labs has highlighted the complexity of the epistemic and affective dynamics of student engagement. Studying labs at Cornell, Sundstrom et al. [83] observed that successful instructor interventions to shift students away from confirmation framing were not always associated with an increase in measures of student agency. And, studying labs at Tufts, Descamps et al. [14] observed students’ confirmation framing was not always associated with a lack of agency; in fact, it seemed to drive productive practices of investigation.
@@
Additionally, Jeon et al. [45] highlight a student’s complex affective experiences of confusion in the Tufts labs: she insisted that she “hates physics,” during side conversation with her lab mates, because their confusing experimental results defied her expectations that physics should align with common sense. At the same time, this students’ feelings of confusion motivated her to take up more epistemic agency during the lab. Thinking in terms of a tradeoff between giving students autonomy and managing their feelings of uncertainty might be a useful heuristic for design decisions, at least within non-traditional labs, Data from study of students’ engagement, however, suggest the effects of different choices are not so clear-cut as “tradeoff” might imply.
@@
As theoretical constructs, framing and agency provide insight into various kinds of design decisions, but they can be taken up in ways that lead to different outcomes. One might focus on decision-making agency [35] or on framing as a phenomenon of an individual’s cognition, but these are far from the only ways to understand and apply framing and agency to pedagogy. Additionally, we cannot ignore the influence of secondary and contextual factors in the divergence of our in-lab guidance structures. At Cornell and UWB, that the instructions provide guidance allows students to do “mini-activities” aimed at building up collaboration skills. The minimal lab instructions at Tufts have the effect of shifting various aspects of classroom management onto the TAs, which is aligned with Hammer’s active interest in TA education, which we take up in the next section.
@@ V.3 Responsive Teaching
We all encourage responsiveness, but, with minimal written guidance, the Tufts labs depend more critically on the instructors’ practices of attending, interpreting, and responding to their students’ thinking. In part, the differences in our expectations of instructors reflect differences in expectations for the materials. Thinking Critically in Physics Labs, the core of the materials used at Cornell and UWB, is designed for wide dissemination and use. It is difficult in general to rely on instructor responsiveness, and so these materials provide more detailed in-lab instructions to guide students toward the intended, rich arenas of experimental inquiry: making useful plots, quantifying uncertainty, refining methods, generating conclusions, and sifting through empirical conflicts.
In contrast, the Tufts labs are only for use at Tufts. That has the advantages and disadvantages of the instructors’ freedom and responsibility to use their judgment as they work with students. We see this as a second “agentive tradeoff” that comes with the more detailed in-lab instructions, analogous to the tradeoff for students: The instructors, too, are guided and constrained with respect to what they may notice and how they may respond.
At UWB, Scherr expects the instructors, who are all faculty, to act with autonomy and agency, including noticing, assessing, and responding to student work during implementation. Scherr’s own modifications to the Thinking Critically in Physics Labs curriculum reflect this aspect of her context. For example, when Scherr introduces the activities around quantitatively comparing datasets, she uses examples from her personal and research interests, while other instructors use different examples. Still, the detailed in-lab instructions function to provide clear representation of pedagogical intentions, while at the same time taking a familiar form, and in these ways may facilitate instructors’ adoption.
@@
At Tufts in contrast, part of the objective for Hammer is TA education and preparation. The TAs are a mix of graduate students and undergraduates, the latter often coming to the position after successful experiences as Learning Assistants [56] in the highly interactive lectures. The minimal written instructions have the effect of shifting responsibility to the TAs for noticing, interpreting, and responding to students. Often this involves helping the students to shift in their framing of lab, such as when students look to the TA for prescriptive guidance over how to proceed.
Hammer’s practice has been to meet with the team of lab TAs for several hours each Friday. The TAs recount snippets of what they noticed in their students’ participation over the week, and they choose examples of students’ written work. They compare interpretations, following practices of educator support for responsive teaching [66], and ideas for how they might respond. They also discuss plans for the coming week, including to suggest refinements to the coming assignment’s wording and design. Experience at Tufts makes it clear that these meetings play an important role in ensuring TAs implement the curriculum as intended and develop skills in responsive teaching.